OpenStudy (anonymous):

Translate to an equation: f $50 poker chips and t $20 poker chips total $400 Graph the equation and use the graph to determine three different combinations of $50 poker chips and $20 poker chips that total $400. Submit your graph through the dropbox

OpenStudy (anonymous):

400=(f x50) + (t x 20)

OpenStudy (anonymous):

\[f 50 + t 20 = 400\] making f = 0. t 20 = 400, t = 20 making t = 0. f50 = 400, f = 8 making f = 4. 4*50 + t20 = 400, t20 = 200, t = 10 Points (f,t) : (0,20) (8,0) (4,10) graph t = 20 - 50/20 f